Two point charges lie on the x-axis: 1.60 μC at x = 1.00 m and –2.30 μC at x = 2.75 m . Part A Find the place(s) where the electric field is zero. Express your answer using two decimal places. If there is more than one answer, separate them by commas.
Answer #1

let us consider that point is to the left of 1.6 uC

at this point electric field due to q1 is equal to electric field due to q2

k q1 / d ^2 = k q2 / ( 1.75 + d )^2

sqrt (q2/q1) = 1.75/d + 1

sqrt ( 2.3/ 1.6) = 1.75/ d +1

d = 8.796 m

x co ordinate of this point = 1 - 8.796 = - 7.796 m

=========

Comment in case any doubt, will reply for sure. Goodluck